OpenStudy (anonymous):

If f(x) = 1 over 2x + 3 and g(x) = −4x − 1, find g[f(4)]. Someone please help me with this!

OpenStudy (anonymous):

First find f(4)

OpenStudy (anonymous):

\[ f(4)=\frac 1{8+3}=\frac 1{11} \]

OpenStudy (anonymous):

Now find \[ g\left( \frac 1 {11} \right)= -4 \left( \frac 1 {11}\right)-1 \]

OpenStudy (anonymous):

So the answer is -1? thank you! I believe I have another I need help with.

OpenStudy (anonymous):

Wait I will find it. Opps.

OpenStudy (anonymous):

No it is not -1

OpenStudy (anonymous):

I am stuck, I am not that good with fractions.

OpenStudy (anonymous):

Sorry :/

OpenStudy (anonymous):

Do I need to turn -4 into -4 over 1?

OpenStudy (anonymous):

So now I divide -15 and 11? Or do I keep it and solve g(x) = -4x-1?

OpenStudy (anonymous):

\[-4 \left( \frac 1 {11}\right)-1 = - \left( \frac 4 {11}\right)-\frac{11}{11}=-\frac{15}{11} \]

OpenStudy (anonymous):

The answer is \[ -\frac {15}{11}\]

OpenStudy (anonymous):

That is not a option on my quiz.

OpenStudy (anonymous):

Is \[f(x)=\frac{1}{2 x+3}\\ or\\ f(x)= \frac 1{2 x} +3 \]

OpenStudy (anonymous):

I am still confused.

OpenStudy (anonymous):

What was f(x) on your quiz? is it the first option above or the second option?

OpenStudy (anonymous):

The first.

OpenStudy (anonymous):

The answer is the one I gave you. There must be a misprint on your quiz. What are the choices?

OpenStudy (anonymous):

21/ -12/ 19/ -19

OpenStudy (anonymous):

None of them is correct. Sorry. In math 1+1=2

OpenStudy (anonymous):

I think your f(x) is \[ f(x)=\frac 12 x + 3 \] In this case the answer is -21

OpenStudy (anonymous):

f(4) =4/2 +3 =5 g(f(4)) =-4(5) -1=-21

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw